OpenStudy (anonymous):

Show that 1/cosx-cosx=(tanx)(sinx)

OpenStudy (anonymous):

Add the 2 expressions on the LHS and use the identity sin^2 + cos^2 =1

OpenStudy (anonymous):

I'm sorry I still don't understand?

OpenStudy (anonymous):

Can u add the 2 expressions on the left side?

OpenStudy (anonymous):

yeah you get cosx/cosx

OpenStudy (anonymous):

\[sin^2x+\cos^2x=1\]\[\tan^2x+1=\frac{1}{\cos^2x}\]\[\sin x \tan x + \cos x=\frac{1}{\cos x}\]

OpenStudy (anonymous):

I have never done a kind of problem that way before..

OpenStudy (anonymous):

But you fully understand this way?

OpenStudy (anonymous):

no I don't understand it at al